OpenStudy (anonymous):

Find the absolute extrema of the function on the closed interval. h(s)= 2/(s-4) [0, 2] my work so far f'(x)=(s-4)(0)/(s-4)^2 = -2/(s-4)^2 f(0)=0 f(0)=0/0-4 = 0 f(2)=2/2-4 =2/-2 =-1 at max (0,0) (-1, 2) min (0,0) but it was all wrong

OpenStudy (freckles):

Ok so you started by trying to find f' ...and f' is looks good \[f'(x)=\frac{(2)'(s-4)-(s-4)'(2)}{(s-4)^2} = \frac{-2}{(s-4)^2} \]

OpenStudy (freckles):

So the only two numbers to check is 0 and 2

OpenStudy (freckles):

did you plug 0 and 2 into the original function?

OpenStudy (freckles):

\[f(0)=\frac{2}{0-4}=? \text{ and } f(2)=\frac{2}{2-4}=?\]

OpenStudy (freckles):

f(0) doesn't equal 0

OpenStudy (freckles):

but f(2) does equal 2/-2=-1 so you did good there

OpenStudy (freckles):

f(0) doesn't equal 0 for sure because 2 will never be 0

OpenStudy (freckles):

\[f(0)=\frac{2}{0-4}=\frac{2}{-4}=\frac{-1}{2}\]

OpenStudy (freckles):

so you have the two points as absolute extrema (0,-1/2) and (2,-1) One of them is the absolute max and the other is the abs min hint: the one with the highest y-value is absolute max and the one with the lowest y value is the absolute min

OpenStudy (freckles):

So your work wasn't all bad

OpenStudy (anonymous):

ok thank you

OpenStudy (freckles):

Do you have any questions?

OpenStudy (anonymous):

no i figured it out thank you